A sky jumper speeds down a slope and exits the ski track in a horizontal direction with a velocity of 25 m/s where the landing incline falls off with a slope of 33°. Solve the following:

A. What is the half time of flight of the sky jumper?

B. What is the range of the sky jumper when it lands?


1
Expert's answer
2021-10-18T11:12:04-0400

"h=gt^2\/2"


"l=vt"


"h=l\cdot\tan33°\to gt^2/2=vt\cdot\tan33°"


"t=2v\tan33°/g=2\cdot25\cdot\tan33°/9.8=3.31\ (s)"


"t_{1/2}=1.66\ (s)" . Answer


"l=25\cdot3.31=82.75 \ (m)" . Answer